# A try for dark energy in quantum field theory: The vacuum energy of neutrino field

### v2

@@ Abstract
In this paper, we try to interpret the origin of dark energy in quantum field theory (QFT). The vacuum energy of a field in QFT is a quartic-divergence input, and its physical contribution to dark energy is calculated in the UV-free scheme. There is a mass-energy window that the active vacuum energy of neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV is expected. The normal hierarchical spectrum of neutrinos is preferred by naturalness, and the neutrino mass window set by dark energy is about 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV.
@@ I Introduction
Physics thrives on crisis Weinberg:1988cp , and the ultraviolet (UV) problem is one of the keys in the development of modern physics. Historically, two dark clouds hung over the classical physics at the end of the 19th century, and then the UV-catastrophe in the dynamical theory of heat became a crisis. The energy quanta was introduced by Planck Planck:1901tja and the UV-catastrophe was solved, and quantum theory was born. In quantum field theory (QFT), the problem of UV divergences reappears, that is, the loop corrections are often UV divergent in a transition process. It means that the transition amplitude obtained by Feynman rules (the physical input) is not directly equal to the physical result (the physical output). How do we extract a finite physical result from a UV-divergent physical input? The paradigm approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator and then removed by counterterms, and this traditional route is based on equivalent transformation and the mathematical structure is $\infty$-$\infty$ - ∞. This paradigm approach is successful in dealing with logarithmic divergences in the standard model (SM). For power-law divergences, there are UV problems, i.e.,
@@
(b) The non-renormalizable Einstein gravity. For power-law divergences of graviton loops, it requires an infinite number of counterterms when one tries to quantize Einstein gravity.
(c) The cosmological constant (dark energy) problem. The vacuum energy in QFT is a quartic divergence, and its contribution (Planck scale) to dark energy exceeds the observation (a density about $(2.3~{}\mathrm{meV})^{4}$) by 120 orders of magnitude.
How do we deal with the above three UV problems in QFT? The free flow of ideas is important when encountering problems. As pointed by Dirac DiracTheEO , the success of renormalization for logarithmic divergences may be similar to the description of one-electron system with further changes required, and the laws of physics are chosen by nature from possible rules that possess mathematical beauty. Feynman Feynman:1986er noted that mathematical self-consistency and legitimacy are important when dealing with UV divergences. Hence, if we go forward, we might have to consider introducing new ideas and new mathematical techniques into QFT to connect UV divergent inputs to the physical reality. For a new route, the requirement is that it first reproduces the successful renormalization results for logarithmic divergences, and then it can give an explanation to the three UV problems caused by power-law divergences. For UV divergences of loops, assuming the physical contributions being local and contributions from UV regions being insignificant, a method of the UV-free scheme Jia:2023dub is introduced to derive loop corrections without UV divergences in calculations, which can reproduce the successful renormalization results for logarithmic divergences in SM. This method takes a new route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$, i.e.,
@@
and there is no UV divergence after the physical analytic continuation $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$ in the UV-free scheme. To the power-law divergences of loops, the UV-free scheme provides an alternative interpretation to the hierarchy problem of Higgs mass within SM Jia:2023dub , and a try for its application to the quantization of Einstein gravity (graviton loops) was investigated in Ref. Jia:2024lyy . In this paper, another power-law divergence, dark energy, is of our concern. The cosmological constant (dark energy) problem is an open fundamental challenge in physics today Weinberg:1988cp ; SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Peebles:2002gy ; Copeland:2006wr . Since the form of vacuum energy ($\varpropto\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}+m^{2}}$) of a field is very similar to the integral of a loop, we will try it in the UV-free scheme in the following.
@@ II The vacuum energy in QFT
$\rho_{0}^{i}=(-1)^{2j}g_{i}\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\sqrt{k^{2}+m_{i}^{2}}\,,$ (4)
where $j$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom of the field. This integral is a quartic divergence, and the form of it is very similar to the integral of a loop. Here the UV-free scheme Jia:2023dub ; Jia:2024lyy is adopted to evaluate the physical contribution of Eq. (4), which is a physical analytic continuation of the integral containing UV divergences in QFT. The Feynman amplitude $\mathcal{T}_{\mathrm{F}}^{i}=\rho_{0}^{i}$ is the physical input, and the Feynman-like amplitude $\mathcal{T}_{\mathrm{F}}^{i}(\xi)$ here can be obtained with $\sqrt{k^{2}+m_{i}^{2}}$ replaced by $\sqrt{k^{2}\!+m_{i}^{2}\!+\xi}$. The physical transition amplitude $\mathcal{T}_{\mathrm{P}}^{i}$, i.e., the physical output, is
@@
$=$ $\!\Big{[}(-1)^{2j}g_{i}\!\mathrm{\int}\!(d\xi)^{3}\!\!\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{3}{16}\frac{1}{\big{(}\sqrt{k^{2}\!+\!m_{i}^{2}\!+\!\xi}\big{)}^{5}}\Big{]}_{\xi\to 0}\!\!+\!C$
$=$ $(-1)^{2j}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log(m_{i}^{2})+C\,\,.$
The result with $C$ set by a typical energy scale $\mu_{\Lambda}$ is
$\mathcal{T}_{\mathrm{P}}^{i}=(-1)^{2j}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}\,.$ (6)
@@
In the limit $m_{i}\to 0$, the result is zero for a massless field. [^1] For another type UV-divergence caused by the discrete sequence, i.e., the Casimir effect, the Riemann zeta function $\zeta(-1)=-\frac{1}{12}$ can be adopted as the analytic continuation. Here the energy scale $\mu_{\Lambda}$ is considered as a scale describing the character of vacuum energy, like the strong interaction scale $\Lambda_{\mathrm{QCD}}$ and the electroweak scale $v$.
III The ripple model and neutrino field
Figure 1: The vacuum energy in the ripple model.
@@ III The ripple model and neutrino field
Let us consider the contribution of a field to dark energy. The result of Eq. (6) indicates that if dark energy comes from the vacuum energy, then the contributions of heavy fields should be suppressed by some unknown mechanisms. Here a small model is introduced to give a possible description. Supposing that the typical energy scale $\mu_{\Lambda}$ is a watershed energy scale to a field with a mass $m_{i}$, the contribution of vacuum energy to dark energy is active in the case of $m_{i}\lesssim\mu_{\Lambda}$, while the contribution of vacuum energy to dark energy is inert in the case of $m_{i}\gg\mu_{\Lambda}$. This is the ripple model, as depicted in Fig. 1. Considering the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy contributing about 71% of the critical density $3H_{0}^{2}/8\pi G$, the small dark energy density is about $(2.3~{}\mathrm{meV})^{4}$ (1 meV $=10^{-3}$ eV), and $\mu_{\Lambda}$ is roughly estimated to be about $10-100$ meV. Thus, the dark energy density is sensitive to the neutrino field, especially the neutrino mass. [^2] Here is more discussions about possible contributions from neutrinos, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , etc. Considering Eq. (6) and dark energy density, an estimate of a typical neutrino mass $m_{\nu}^{\mathrm{tp}}$ is obtained, with $m_{\nu}^{\mathrm{tp}}\sim 2.3\times(32\pi^{2})^{\frac{1}{4}}\approx$ 9.7 meV (here $|\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}|^{\frac{1}{4}}\sim 1$ is adopted). The vacuum energy of a neutrino field with a mass about $m_{\nu}^{\mathrm{tp}}$ can naturally produces the required dark energy density. As a rough prediction, there should be a neutrino with a mass $\sim$ 10 meV.
@@
Figure 2: The normal hierarchical spectra of neutrinos (normal ordering) with dark energy density considered. The solid curves are masses of neutrinos with the mass of $\nu_{1}$ as the input, and the bands with a 5% change in neutrino mass are plotted for a convenient naturalness estimation. The dotted curves are the values of $\mu_{\Lambda}$ required by the dark energy density $(2.3~{}\mathrm{meV})^{4}$ with neutrino masses inputted. The region between the two dashed lines is the neutrino mass window.
@@
Figure 3: The inverted hierarchical spectra of neutrinos (inverted ordering) with dark energy density considered. The solid curves are masses of neutrinos with the mass of $\nu_{3}$ as the input. The dotted curve is the values of $\mu_{\Lambda}$ required by the dark energy density $(2.3~{}\mathrm{meV})^{4}$.
@@
Now we turn to neutrino masses. The masses of three neutrinos are $m_{1}$, $m_{2}$ and $m_{3}$, and the mass differences between them can be extracted by neutrino oscillations. The mass-splitting results from the Particle Data Group ParticleDataGroup:2024cfk are $\Delta m_{21}^{2}\approx 7.41\times 10^{-5}\mathrm{eV}^{2}$, $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx 2.437\times 10^{-3}\mathrm{eV}^{2}$ (normal ordering) and $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx-2.498\times 10^{-3}\mathrm{eV}^{2}$ (inverted ordering). Considering the naturalness with an effective mass $m_{\nu}^{\mathrm{tp}}$, the quasidegenerate spectrum with $m_{1}\simeq m_{2}\simeq m_{3}\gg\sqrt{|\Delta m_{32}^{2}|}$ is not favored. It is possible to give an estimate on the mass spectrum of neutrinos with $m_{\nu}^{\mathrm{tp}}$, and the results obtained are model-dependent. Here we consider a specific implementation of the ripple model by a test factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ describing the active contribution, and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$. For a field with a mass $m_{i}\gg\mu_{\Lambda}$, its contribution to the dark energy density is negligible due to the tiny value of $g_{i}^{\ast}$ (this is similar to the freezing of effective degrees of freedom of heavy particles at a low temperature in the early universe). In this case, the observation of dark energy density indicates that neutrino fields play a major role.
The hierarchical spectra of neutrinos for normal ordering and inverted ordering with dark energy density considered are shown in Fig. 2 and Fig. 3, respectively. In Fig. 3, considering the naturalness of $\mu_{\Lambda}$, it can be seen that the inverted hierarchical spectrum is not favored. Let us look at the normal ordering result shown in Fig. 2. There is a region that the neutrino masses at given $\mu_{\Lambda}$ can naturally produce the active vacuum energy required for the dark energy density, i.e., a mass-energy window. Likewise, the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV, and the total neutrino mass is 67.5 meV $\lesssim m_{1}+m_{2}+m_{3}\lesssim$ 87.4 meV. This mass window can be tested by the future experiments.
@@
Here we give a brief discussion about the Hubble tension. A slightly higher $H_{0}=(73.04\pm 1.04$) km s-1 Mpc-1 Riess:2016jrr ; Riess:2021jrx from the nearby universe and a slightly lower $H_{0}=(67.4\pm 0.5$) km s-1 Mpc-1 Planck:2018vyg from the early universe indicate a Hubble tension. For $H_{0}\approx$ 67.4 km s-1 Mpc-1 and $H_{0}\approx$ 73 km s-1 Mpc-1, the corresponding dark energy densities are about $(2.24~{}\mathrm{meV})^{4}$ and $(2.37~{}\mathrm{meV})^{4}$, respectively. If the energy scale $\mu_{\Lambda}$ has a slow-running behavior and traverses the upper part of the mass-energy window in Fig. 2, the Hubble tension can be explained by this slow-running behavior. If the energy scale $\mu_{\Lambda}$ is fixed, new explanations and mechanisms need to be taken into account.

### v4

@@ Abstract
The quartic-divergent vacuum energy poses an ultraviolet (UV) challenge (the cosmological constant problem) in probing the nature of dark energy. Here we try to evaluate the contribution of the vacuum energy to dark energy with a method of the UV-free scheme. The result indicates that it is not a problem in the UV region but a question of the contributions of heavy fields being suppressed. Then, we propose an effective description via scale decoupling. The parameter spaces suggest that the active vacuum energy of neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV is expected. The normal ordering neutrinos are preferred by naturalness, and the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV.
@@ I Introduction
Physics thrives on crisis Weinberg:1988cp , and the ultraviolet (UV) problem is one of the keys in the development of modern physics. In the context of the UV catastrophe in classical physics, the energy quanta were introduced by Planck Planck:1901tja while studying blackbody radiation problems, which led to the birth of quantum theory. In quantum field theory (QFT), the problem of UV divergences reappears, that is, loop corrections of a transition process are often UV divergent. It means that the transition amplitude obtained by Feynman rules (the physical input) is not directly equal to the physical result (the physical output). To extract the finite physical result from a UV-divergent input, the traditional approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator based on equivalent transformation and then removed by counterterms, with the mathematical structure of this route being $\infty$-$\infty$ - ∞. This traditional approach is successful in dealing with logarithmic divergences in the standard model (SM). When we leave the logarithmic region and move on, there are UV problems for power-law divergences, i.e.,
@@
(b) The non-renormalizable Einstein gravity. When one tries to quantize Einstein gravity, an infinite number of counterterms are required for graviton loops exhibiting power-law divergences.
(c) The cosmological constant problem. The vacuum energy density in QFT is a quartic divergence, and its contribution (Planck scale) exceeds the critical density for a nearly flat universe by 120 orders of magnitude, with the dark energy density $\sim(2.3~{}\mathrm{meV})^{4}$.
How do we deal with the above three UV problems? The traditional approach exhibits UV limitations in handling power-law divergences. In this case, the free flow of ideas is important. Dirac DiracTheEO viewed renormalization as a mathematical framework requiring refinement to address UV divergences; the physical laws chosen by nature were found to possess mathematical beauty. Feynman Feynman:1986er emphasized that mathematical self-consistency and rigor are essential in addressing UV divergences. In the vision of Dyson Dyson:1949ha and Schwinger Schwinger , future approaches to loop corrections would be based on finite quantities. Wilsonian effective theory Wilson:1971bg decouples scales by encoding UV effects into low-energy parameters via a physical cutoff, enabling systematic organization of loop contributions. Hence, if we go forward, we might have to consider introducing new physical ideas and new mathematical techniques to connect UV divergent inputs to the physical reality. For a new route, the requirement is that it first reproduces the successful renormalization results for logarithmic divergences, and then it can give an interpretation to the three UV problems caused by power-law divergences. For UV divergences of loops, assuming the physical contributions being local and insensitive to UV regions, a method of the UV-free scheme Jia:2023dub is introduced to derive loop corrections, which can reproduce the renormalization results for logarithmic divergences in SM. This method takes a route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$,
@@
and UV divergences are systematically eliminated through $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$ without counterterms. Loop corrections in this method are constructed from local finite quantities that maintain scale decoupling, while ensuring UV-safe loop contributions through locality-based analytic continuation. To power-law divergences of loops, this method provides an interpretation to the hierarchy problem of Higgs mass within SM Jia:2023dub , and a try for its application to Einstein gravity (graviton loops) was investigated in Ref. Jia:2024lyy . Here another power-law divergence, the vacuum energy, is of our concern. The cosmological constant (dark energy) problem is an open fundamental challenge in physics today Weinberg:1988cp ; SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Peebles:2002gy ; Copeland:2006wr . In exploring the nature of dark energy, an unavoidable question is the contribution of the vacuum energy density $\rho_{\mathrm{vac}}$ that corresponds to a cosmological constant ($\Lambda=8\pi G\rho_{\mathrm{vac}}$). Since the form of the vacuum energy of a field ($\varpropto\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}+m^{2}}$) is similar to the integral of a loop, we will try it in the UV-free scheme in the following.
@@ II The vacuum energy in QFT
$\rho_{0}^{i}=(-1)^{2j_{i}}g_{i}\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\sqrt{k^{2}+m_{i}^{2}}\,,$ (4)
where $j_{i}$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom of the field. This integral is a quartic divergence. Here the UV-free scheme Jia:2023dub ; Jia:2024lyy is adopted to evaluate the physical contribution of Eq. (4), which is a physical analytic continuation of the integral containing UV divergences. The Feynman amplitude $\mathcal{T}_{\mathrm{F}}^{i}=\rho_{0}^{i}$ is the physical input, and the Feynman-like amplitude $\mathcal{T}_{\mathrm{F}}^{i}(\xi)$ here can be obtained with $\sqrt{k^{2}+m_{i}^{2}}$ replaced by $\sqrt{k^{2}\!+m_{i}^{2}\!+\xi}$. The physical transition amplitude $\mathcal{T}_{\mathrm{P}}^{i}$, i.e., the physical output, is
@@
$=$ $\!\Big{[}(-1)^{2j_{i}}g_{i}\!\mathrm{\int}\!(d\xi)^{3}\!\!\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{3}{16}\frac{1}{\big{(}\sqrt{k^{2}\!+\!m_{i}^{2}\!+\!\xi}\big{)}^{5}}\Big{]}_{\xi\to 0}\!\!+\!C$
$=$ $(-1)^{2j_{i}}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log(m_{i}^{2})+C\,\,.$
The result with $C$ set by a reference energy scale $\mu_{\Lambda}$ is
$\mathcal{T}_{\mathrm{P}}^{i}=(-1)^{2j_{i}}g_{i}\frac{1}{64\pi^{2}}m_{i}^{4}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}\,.$ (6)
@@
In the limit $m_{i}\to 0$, the result is zero for a massless field. [^1] For another type UV-divergence caused by the discrete sequence, i.e., the Casimir effect, the Riemann zeta function $\zeta(-1)=-\frac{1}{12}$ can be adopted as the analytic continuation. The energy scale $\mu_{\Lambda}$ is considered as a scale describing the character of vacuum energy, like the strong interaction scale $\Lambda_{\mathrm{QCD}}$ and the electroweak scale $v$.
Here is an interpretation of scales in QFT. The physical result of total tree-level and loop-level contributions in a scattering process is independent of the choice of an energy scale $\mu$, and the total physical result can be equivalently described by a tree-level process with an effective coupling at an effective scale $\mu_{\mathrm{eff}}$. The energy scale $\mu_{\Lambda}$ of the vacuum energy is similar to the scale $\mu_{\mathrm{eff}}$ in the total physical result, which is different from the free parameter $\mu$ in order-by-order expansions. In addition, $\mu_{\mathrm{eff}}$ changes with the scattering energy, while the scale $\mu_{\Lambda}$ describes the reference scale for the ground states of all fields, i.e., all zero-point fields have the same reference scale $\mu_{\Lambda}$. Moreover, $\mu_{\Lambda}$ may be a fixed or slow-running scale with the expansion of the universe.
III The ripple description and neutrino field
Figure 1: The vacuum energy in the ripple description.
@@ III The ripple description and neutrino field
Let us consider the contribution of a field to dark energy. The result of Eq. (6) indicates that it is no longer a problem in the UV region. The actual question is revealed, that is, the contributions of heavy fields should be suppressed by some unknown mechanisms that are beyond the current understanding of QFT. If we analyze the question with Wilsonian framework Wilson:1971bg , to vacuum energy, there may be scale decoupling between microscopic quantity and macroscopic effect. We will try to give it an effective description in the view of scale decoupling, that is, the total zero-point energy is the microscopic energy foundation as a background, and quantum fluctuations with cumulative contributions characterized by $\mu_{\Lambda}$ manifest macroscopic effects through fluctuation-carrying fields. The scale $\mu_{\Lambda}$ is a watershed scale to a carrier field with a mass $m_{i}$, that is, the contribution of a field’s vacuum energy to dark energy is active, inert in the case of $m_{i}\lesssim\mu_{\Lambda}$, $m_{i}\gg\mu_{\Lambda}$, respectively. This picture likes ripples on a lake, necessitating $\mu_{\Lambda}-$characterized fluctuation factor for ground states of fields (e.g., the energy quanta distribution in blackbody radiation research). This is the ripple description, as depicted in Fig. 1. Considering the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy contributing 71% of the critical density $3H_{0}^{2}/8\pi G$, the dark energy density is about $(2.3~{}\mathrm{meV})^{4}$ (1 meV $=10^{-3}$ eV), and $\mu_{\Lambda}$ is roughly estimated to be $10-100$ meV. Thus, the dark energy density is sensitive to the neutrino field, especially the neutrino mass. [^2] Here is more discussions about possible neutrino contributions, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , the cosmological quantum vacuum energy of axion and neutrino deVega:2023fvw , etc. Considering Eq. (6) and dark energy density, a typical neutrino mass is $m_{\nu}^{\mathrm{tp}}\sim 2.3\times(32\pi^{2})^{\frac{1}{4}}\approx$ 9.7 meV (here $|\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}|^{\frac{1}{4}}\sim 1$). As a rough prediction, to produce the required dark energy density, there should be a neutrino with a mass $\sim$ 10 meV.
@@
Figure 2: The normal ordering neutrinos with the dark energy density $(2.3~{}\mathrm{meV})^{4}$. The solid curves are masses of neutrinos with the mass of $\nu_{1}$ as the input, and the bands with a 5% change in neutrino mass are plotted for a convenient naturalness estimation. The dotted curves are the values of $\mu_{\Lambda}$ required by the dark energy density with neutrino masses inputted. The region between the two dashed lines is the neutrino mass window.
@@
Figure 3: The inverted ordering neutrinos with the dark energy density $(2.3~{}\mathrm{meV})^{4}$. The solid curves are masses of neutrinos with the mass of $\nu_{3}$ as the input. The dotted curve is the values of $\mu_{\Lambda}$ required by the dark energy density.
@@
Now we turn to neutrino masses. The masses of three neutrinos are $m_{1}$, $m_{2}$ and $m_{3}$, and the mass differences between them can be extracted by neutrino oscillations. The mass-splitting results ParticleDataGroup:2024cfk are $\Delta m_{21}^{2}\approx 7.41\times 10^{-5}\mathrm{eV}^{2}$, $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx 2.437\times 10^{-3}\mathrm{eV}^{2}$ (normal ordering) and $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx-2.498\times 10^{-3}\mathrm{eV}^{2}$ (inverted ordering). Considering the naturalness with an effective mass $m_{\nu}^{\mathrm{tp}}$, the quasidegenerate spectrum with $m_{1}\simeq m_{2}\simeq m_{3}\gg\sqrt{|\Delta m_{32}^{2}|}$ is not favored. It is possible to give an estimate on the mass spectrum of neutrinos. Here we consider a specific implementation of the ripple description by a test fluctuation factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ in Gaussian distribution to characterize the active contribution of a field (another possible form $e^{-m_{i}/\mu_{\Lambda}}$ in Boltzmann distribution is not preferred by the naturalness of parameter spaces), and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ ($\rho_{\mathrm{vac}}=\sum_{i}(-1)^{2j_{i}}g_{i}^{\ast}\frac{m_{i}^{4}}{64\pi^{2}}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}$). For a field with a mass $m_{i}\gg\mu_{\Lambda}$, its contribution to the dark energy density is negligible due to the tiny value of $g_{i}^{\ast}$ (this is similar to the freezing of effective degrees of freedom of heavy particles at a low temperature in the early universe). The hierarchical spectra of neutrinos for normal ordering and inverted ordering with the required dark energy density are shown in Fig. 2 and Fig. 3, respectively. In Fig. 3, considering the naturalness of $\mu_{\Lambda}$, the inverted ordering is not favored. For the normal ordering shown in Fig. 2, there is a region that neutrino masses at given $\mu_{\Lambda}$ can naturally produce the active vacuum energy required by the dark energy density, i.e., the parameter spaces forming a mass-energy window. Likewise, the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV, and the total neutrino mass is 67.5 meV $\lesssim m_{1}+m_{2}+m_{3}\lesssim$ 87.4 meV. This mass window can be tested by the future experiments.
@@
Let us give a brief discussion about the Hubble tension. A slightly higher $H_{0}=(73.04\pm 1.04$) km s-1 Mpc-1 Riess:2016jrr ; Riess:2021jrx from the nearby universe and a slightly lower $H_{0}=(67.4\pm 0.5$) km s-1 Mpc-1 Planck:2018vyg from the early universe indicate a possible Hubble tension (the tension may be alleviated Freedman:2024eph ). For $H_{0}\approx$ 67.4 km s-1 Mpc-1 and $H_{0}\approx$ 73 km s-1 Mpc-1, the corresponding dark energy densities are about $(2.24~{}\mathrm{meV})^{4}$ and $(2.37~{}\mathrm{meV})^{4}$, respectively. If the energy scale $\mu_{\Lambda}$ has a slow-running behavior and traverses the upper part of the mass-energy window in Fig. 2, the Hubble tension can be explained by this slow-running behavior, i.e., a running cosmological constant. If $\mu_{\Lambda}$ is fixed, new explanations and mechanisms need to be taken into account.